# Efficient Egocentric Human Activity Recognition

### v1

@@ 4.3.1 IMU-based Human Activity Recognition
EgoExo4D MMEA Mantis 92.25 98.10 99.16
@@
Table 3: Cross-Dataset Generalization results showing transferability to unseen data. $\dagger$ indicates the model was supervised fine-tuned on the train dataset.
@@
Acc@1 Acc@3 Acc@5 Acc@1 Acc@3 Acc@5 Acc@1 Acc@3 Acc@5
MOMENTS VideoMAEB-k400 IMU2CLIP 41.39 65.82 76.45 69.77 94.61 98.81 78.42 91.19 94.60
@@
MOMENTS TimeSformerB-k400 IMU2CLIP 43.59 68.17 78.06 70.73 95.15 98.98 78.80 92.17 95.14
@@
Mantis VideoMAEB-k400 IMU2CLIP 54.43 75.92 83.89 79.09 97.57 99.44 90.88 98.63 99.47
@@
Mantis TimeSformerB-k400 IMU2CLIP 54.83 76.04 84.03 79.46 97.68 99.44 91.41 98.71 99.47
@@
Table 4: Result for different video teacher IMU student model combinations on Ego4D, EgoExo4D, and MMEA.
4.3.2 Cross-Dataset Generalization
The goal of this task is to measure the capability of a model to generalize its learned representations to other datasets that differ significantly from its training set. Strong cross-dataset generalization is critical for deploying HAR models in diverse and realistic environments, where variations in sensor setups, activities, and user behaviors are common. To evaluate this capability, we train models on one dataset and test them on another without additional fine-tuning, following the evaluation protocol in Section 3.2 to assess zero-shot generalization. We experiment with the following setups: (1) EgoExo4D $\rightarrow$ Ego4D, (2) Ego4D $\rightarrow$ EgoExo4D, (3) Ego4D $\rightarrow$ MMEA, and (4) EgoExo4D $\rightarrow$ MMEA.
Results. Table 3 presents the results of our cross-dataset generalization experiments. Notably, COMODO improves over IMU2CLIP and supervised fine-tuned Mantis baselines in all cases. In the EgoExo4D → Ego4D setup, COMODO outperforms IMU2CLIP by 4.87% in Acc@1 while also surpassing the supervised fine-tuned Mantis by 4.04%. When transferring from Ego4D to EgoExo4D, COMODO achieves the highest Acc@1 of 82.54%, improving upon the Mantis baseline by 1.75%. On the MMEA dataset, COMODO maintains its superiority, surpassing Mantis by 2.5% and 2.58% when trained on Ego4D and EgoExo4D, respectively. These results underscore COMODO’s robustness in adapting to datasets with significantly different distributions, further demonstrating the effectiveness of cross-modal self-supervised knowledge transfer in enhancing IMU representation learning.
4.4 Ablation Study
4.4.1 Different Teacher & Student Networks
To evaluate the generalizability of COMODO across different models, we conduct experiments using two IMU student models (Mantis [11] and MOMENT-small [16]) and two video teacher models (VideoMAE-Base-k400222https://huggingface.co/MCG-NJU/videomae-base-finetuned-kinetics [40] and TimeSformer-Base-k400 [2]). The results in Table 4 demonstrate that COMODO consistently improves performance across the vast majority of teacher-student configurations, validating its robustness in different settings.
We observe that TimeSformer-Base-k400 generally outperforms VideoMAE-Base-k400, suggesting that a stronger spatio-temporal feature extractor benefits IMU representation learning. Similarly, Mantis surpasses MOMENT-small as an IMU student, aligning with its superior standalone performance in Table 1. Notably, COMODO achieves significant gains across all networks, demonstrating its teacher-student model agnosticism and adaptability to various video and time-series models.
Figure 3: Impact of queue size on accuracy across datasets.
4.4.2 FIFO Queue Size
@@
The optimal queue size $|Q|$ is dataset-specific, as we observe that different datasets benefit from different queue configurations. Our ablation study (Figure 3) reveals that increasing queue size does not always yield better results. Instead, the relationship between queue size and performance is influenced by the characteristics of the dataset:
@@
Ego4D ($|Q|$ = 16384): As a large-scale dataset with diverse activities (31 classes), increasing the queue size generally improves performance by providing a richer set of negative samples for contrastive learning. However, we observed that the largest queue size tested (65536) does not yield the best performance. Prior research [29] also suggests that the optimal queue size is not necessarily the largest available, and an overly large queue may not always lead to improved results.
@@
EgoExo4D ($|Q|$ = 1024): Empirical results indicate that increasing the queue size initially improves performance, but beyond $|Q|$ = 1024, further increasing the queue size leads to a decline in accuracy (Figure 3). Unlike Ego4D and MMEA, EgoExo4D has only 8 activity classes, leading to highly concentrated feature distributions. When the queue size is large, excessive redundant negatives from the same limited set of categories create sharp similarity spikes in the distribution. This makes it harder for the student model to align with the teacher’s representation. Prior work [20] suggests that KL divergence is easier to minimize when the distribution has a single dominant mode, whereas in the distributions of multiple distinct modes, forward KL forces the model to cover all modes, potentially assigning probability mass to less relevant regions. A moderate queue size helps regulate the number of distinct similarity peaks, stabilizing training and improving teacher-student alignment.
@@
Train Method Acc@1 Acc@3 Acc@5
@@
Table 5: Comparison of mean pooling and concat pooling for IMU representation learning across different datasets.
4.4.3 Pooling Methods
@@
Table 5 shows that concat pooling consistently outperforms mean pooling across all datasets. On Ego4D, it improves Acc@1 by 1.82%, while on EgoExo4D, the gain is 0.98%. The most significant improvement is observed on MMEA, where Acc@1 increases by 6.84%, indicating that maintaining per-channel information is especially beneficial for this dataset. These results suggest that mean pooling overly compresses features, leading to information loss, while concat pooling retains richer representations, enhancing classification accuracy. Given its superior performance across all datasets, we adopt concat pooling as the default method in our framework, as discussed in Section 4.2.
Figure 4: Accuracy of distillation methods across datasets.
4.4.4 Distillation Methods
To assess the effectiveness of our distillation strategy, we compare our strategy against two other distillation methods.
@@ 4.4.4 Distillation Methods
(1) L2 Distance Minimization [36]: L2 distance minimization directly aligns the student and teacher representations by minimizing the squared L2 distance between the IMU student embedding $z_{xi}$ and the video teacher embedding $z_{vi}$ for an identical paired video-IMU sample ($v_{i},x_{i}$). This method assumes that the feature spaces of different modalities are inherently aligned. As a result, L2 minimization can lead to suboptimal feature alignment.
@@
(2) InfoNCE-based Contrastive Learning: InfoNCE optimizes the student encoder by maximizing the similarity between positive video-IMU pairs while minimizing similarity with randomly sampled negative instances. While effective in learning discriminative features, this method treats all negative samples equally, potentially misclassifying semantically similar negatives as incorrect, which limits its adaptability in cross-modal distillation.
@@
(3) COMODO (ours): COMODO extends contrastive learning by distilling the similarity distribution rather than relying solely on instance-wise alignment. Unlike InfoNCE, which assumes all negatives are equally dissimilar, COMODO preserves the structure of the teacher’s similarity space, leading to more robust and generalizable feature representations. Figure 4 shows that COMODO consistently outperforms L2 and InfoNCE across all datasets, demonstrating the effectiveness of similarity distribution distillation in cross-modal representation learning.
5 Conclusion
In this work, we introduced COMODO, a cross-modal self-supervised distillation framework that leverages pretrained video representations to enhance IMU-based egocentric HAR. By distilling rich semantic knowledge from a frozen video encoder to an IMU encoder through a dynamic instance queue, COMODO effectively mitigates the challenges posed by limited sensor data and costly annotations. Extensive experiments across multiple egocentric HAR benchmarks demonstrate that COMODO achieves results comparable to or even surpassing fully supervised baselines. Notably, our method exhibits strong cross-dataset generalization, underscoring its robustness and practical utility. COMODO’s adaptability to various pretrained video and sensor models offers a viable framework for further investigating cross-modal knowledge transfer from visual models to resource-constrained sensors, enhancing future research in visually-guided, sensor-based HAR.
6 Acknowledgment

### v2

@@ Results.
EgoExo4D MMEA CrossHAR 79.86 94.45 97.72
Mantis 92.25 98.10 99.16
@@
Table 4. Cross-Dataset Generalization results showing the method’s ability to transfer to completely unseen datasets, where both the activity classes and devices differ from those in the training set. $\dagger$ indicates supervised fine-tuning on the source training dataset.
4.4. Cross-Dataset Generalization
The ultimate goal of representation learning is to learn features that are not only effective on the training domain but also generalize to new, unseen environments. Strong cross-dataset generalization is critical for deploying HAR models in diverse and realistic environments, where variations in sensor setups, activities, and user behaviors are common. To measure this capability, we designed a rigorous cross-dataset evaluation task. This task assesses how well the representations learned on one dataset (the source dataset) can be utilized on a completely different dataset (the target dataset) without additional fine-tuning, following the evaluation protocol in Section 3.2 to assess out-of-domain generalization capability of the learned encoder. We experiment with the following setups: (1) EgoExo4D $\rightarrow$ Ego4D, (2) Ego4D $\rightarrow$ EgoExo4D, (3) Ego4D $\rightarrow$ MMEA, and (4) EgoExo4D $\rightarrow$ MMEA.
Results.
Table 4 presents the results of our cross-dataset generalization experiments. Notably, COMODO improves over IMU2CLIP, CrossHAR and supervised fine-tuned Mantis baselines in all cases. In the EgoExo4D → Ego4D setup, COMODO outperforms IMU2CLIP by 4.87% in A@1 while also surpassing the supervised fine-tuned Mantis by 4.04%. Furthermore, COMODO demonstrates a substantial performance lead over CrossHAR (+26.07% A@1), which suggests that intra-modal self-supervision struggles to bridge such large domain gaps without external semantic guidance. When transferring from Ego4D to EgoExo4D, COMODO achieves the highest A@1 of 82.54%, improving upon Mantis, IMU2CLIP, and CrossHAR by 1.75%, 2.84%, and 14.32%. On the MMEA dataset, COMODO maintains its superiority, surpassing Mantis by 2.5% and 2.58% when trained on Ego4D and EgoExo4D, respectively. The findings from this cross-dataset evaluation are highly significant for practical HAR applications. They underscore COMODO is not merely improving performance on a specific data distribution and setting, but is fundamentally enhancing the representation quality and transferability of the IMU encoder, further demonstrating the effectiveness of cross-modal self-supervised knowledge transfer in enhancing IMU representation learning.
4.5. Efficiency Analysis
Inference Efficiency and Deployability
To validate the deployability of COMODO, we evaluate its inference efficiency in a single-threaded CPU setting on an AMD EPYC 9254 processor, simulating a resource-constrained environment. As illustrated in Figure 6, COMODO achieves a superior balance between recognition accuracy and computational cost. All latency measurements are reported as the average per-sample inference time over 100 samples, and FLOPs are computed for a single forward pass.
Figure 6. Comparison of performance and computational efficiency. We report mean A@1 accuracy across three datasets versus FLOPs (left) and inference latency (right), both in log scale. Bubble size denotes model size in terms of parameters. The red dashed line marks the 5-second input window.
[…]
Furthermore, compared to IMU-based models like DeepConvLSTM and Attend and Discriminate, COMODO delivers a substantial accuracy gain (up to +18.2% mean A@1) while maintaining millisecond-level inference latency suitable for deployment. Moreover, despite sharing the same architecture and hardware footprint as IMU2CLIP, COMODO consistently achieves higher accuracy (+3.6% mean A@1), demonstrating that similarity distribution distillation enhances semantic representations without introducing additional inference overhead.
Training Computational Cost and Scalability
In addition to on-device inference efficiency, we also evaluate the training-time computational cost, which is critical for scalability on large datasets. Since COMODO relies on a frozen video teacher and only updates the lightweight IMU encoder, the training process is efficient. On a single NVIDIA L40S GPU, the average training time per epoch is 232.3s for Ego4D, 98.95s for EgoExo4D, and 11.8s for MMEA. Moreover, COMODO employs an offline feature extraction strategy (see Section 4.2), where video embeddings are pre-computed only once for each dataset and cached, and directly reused during IMU encoder training without invoking video inference. This design decouples the training-time computational cost from the complexity of the video teacher, enabling efficient training even when leveraging high-capacity video models. This low computational barrier ensures that COMODO is easily scalable to large datasets without requiring prohibitive computational resources.
4.6. Ablation Study
4.6.1. Different Teacher & Student Networks.
To evaluate the generalizability of COMODO across different models, we conduct experiments using two IMU student models (Mantis  (Feofanov et al., 2025) and MOMENT-small  (Goswami et al., 2024)) and two video teacher models (VideoMAE-Base-k400 (Tong et al., 2022) and TimeSformer-Base-k400  (Bertasius et al., 2021)). The results in Table 5 reveal that COMODO consistently and substantially outperforms the IMU2CLIP across every teacher-student configuration on all datasets.
@@ 4.6. Ablation Study
A@1 A@3 A@5 A@1 A@3 A@5 A@1 A@3 A@5
MOMENTS VideoMAEB IMU2CLIP 41.39 65.82 76.45 69.77 94.61 98.81 78.42 91.19 94.60
@@
MOMENTS TimeSformerB IMU2CLIP 43.59 68.17 78.06 70.73 95.15 98.98 78.80 92.17 95.14
@@
Mantis VideoMAEB IMU2CLIP 54.43 75.92 83.89 79.09 97.57 99.44 90.88 98.63 99.47
@@
Mantis TimeSformerB IMU2CLIP 54.83 76.04 84.03 79.46 97.68 99.44 91.41 98.71 99.47
@@
Table 5. Result for different video teacher-IMU student model combinations on Ego4D, EgoExo4D, and MMEA. The best are indicated in bold, and the second-best are underlined.
We observe two key trends. First, regarding the student models, Mantis consistently proves to be a more capable learner than MOMENT-small, achieving significantly higher performance regardless of the teacher. This aligns with its superior standalone results in Table 3. Second, regarding the video teacher, we observe that both VideoMAE-Base-k400 and TimeSformer-Base-k400 act as highly effective knowledge sources. While neither demonstrates universal superiority across all settings, their standalone performances are strong and highly comparable. This highlights a crucial strength of our method: COMODO’s effectiveness is not contingent on a specific teacher architecture, demonstrating strong model-agnostic capabilities and adaptability to both video and time-series models. This is further evidenced by the fact that the combination of the stronger Mantis student with either teacher consistently yields the best and second-best results overall.
Video Teacher Performance
To contextualize the performance of our IMU-only model, we present the capabilities of the video teacher models in Table 6. This serves as a proxy for the intrinsic difficulty of each classification task. We report two sets of teacher results: pre-trained models (VideoMAEB-k400 and TimeSformerB-k400) evaluated by training an SVM classifier on their extracted features, which aligns with the evaluation setting for our student model in the main paper. Additionally, we include results for these same models after being supervised fine-tuned on each training dataset $\dagger$}, evaluated using their classifier heads setting.
Methods Ego4D EgoExo4D MMEA
A@1 A@3 A@5 A@1 A@3 A@5 A@1 A@3 A@5
COMODO (Ours) 59.13 78.79 85.65 84.92 98.28 99.59 92.48 99.01 99.77
VideoMAEB-k400 74.31 87.80 92.15 99.27 99.90 100 92.78 99.32 99.62
TimeSformerB-k400 76.16 88.99 92.43 99.63 99.96 100 93.47 99.01 99.54
[…]
Table 6. Performance comparison between the IMU student (COMODO) and its video teachers. $\dagger$} denotes supervised finetuned on the training dataset.
The analysis reveals that the video teachers achieve near-perfect accuracy on EgoExo4D and MMEA, signifying that the activities in these datasets are highly distinguishable from the visual modality. In contrast, the lower scores on Ego4D confirm its greater complexity. Impressively, our IMU-only COMODO model successfully closes a significant portion of the modality gap, especially on the more challenging Ego4D dataset. On EgoExo4D and MMEA, COMODO’s performance is remarkably close to that of its video teachers, demonstrating that our distillation framework effectively transfers rich semantic knowledge from the powerful video modality to the far more lightweight IMU modality in a self-supervised way.
4.6.2. FIFO Queue Size
@@
The optimal queue size $|Q|$ is dataset-specific, as different datasets benefit from different queue configurations. Our ablation study (Figure 7) reveals that increasing queue size does not always yield better results. Instead, the relationship between queue size and performance is influenced by the characteristics of the dataset:
Figure 7. Impact of queue size on accuracy across datasets.
@@
Ego4D ($|Q|$ = 16384): As a large-scale dataset with diverse activities (31 classes), increasing the queue size generally improves performance by providing a richer set of negative samples for contrastive learning. However, we observed that the largest queue size tested (65536) does not yield the best performance. Prior research  (Limkonchotiwat et al., 2022) also suggests that the optimal queue size is not necessarily the largest, and an overly large queue may not always lead to improved results.
@@
EgoExo4D ($|Q|$ = 1024): Empirical results indicate that increasing the queue size initially improves performance, but beyond $|Q|$ = 1024, further increasing the queue size leads to a decline in accuracy (Figure 7). Unlike Ego4D and MMEA, EgoExo4D has only 8 activity classes, leading to highly concentrated feature distributions. When the queue size is large, excessive redundant negatives from the same limited set of categories create sharp similarity spikes in the distribution. This makes it harder for the student model to align with the teacher’s representation. Prior work  (Gu et al., 2024) suggests that KL divergence is easier to minimize when the distribution has a single dominant mode, whereas in the distributions of multiple distinct modes, forward KL forces the model to cover all modes, potentially assigning probability mass to less relevant regions. A moderate queue size helps regulate the number of distinct similarity peaks, stabilizing training and improving teacher-student alignment.
@@
4.6.3. Pooling Methods
Train Method A@1 A@3 A@5
@@ 4.6.3. Pooling Methods
Table 7. Comparison of mean pooling and concat pooling for IMU representation learning across different datasets.
@@
Table 7 shows that concat pooling consistently outperforms mean pooling across all datasets. On Ego4D, it improves A@1 by 1.82%, while on EgoExo4D, the gain is 0.98%. The most significant improvement is observed on MMEA, where A@1 increases by 6.84%, indicating that maintaining per-channel information is especially beneficial for this dataset. These results suggest that mean pooling overly compresses features, leading to information loss, while concat pooling retains richer representations, enhancing classification accuracy. Given its superior performance across all datasets, we adopt concat pooling as the default method in our framework, as discussed in Section 4.2.
4.6.4. Distillation Methods.
To empirically validate the theoretical analysis presented in Section 3.1.1, we compare our strategy against two representative baseline methods.
@@
(1) L2 Distance Minimization  (Romero et al., 2015): L2 distance minimization directly aligns the student and teacher representations by minimizing the squared L2 distance between the IMU student embedding $z_{xi}$ and the video teacher embedding $z_{vi}$ for an identical paired video-IMU sample ($v_{i},x_{i}$). This method assumes that the feature spaces of different modalities are inherently aligned. As a result, L2 minimization can lead to suboptimal feature alignment.
@@
(2) InfoNCE-based Contrastive Learning: InfoNCE optimizes the student encoder by maximizing the similarity between positive video-IMU pairs while minimizing similarity with randomly sampled negative instances. As discussed in Section 3.1.1, this formulation corresponds to the “hard” contrastive objective, which imposes a one-hot target distribution. While effective in learning discriminative features, this method treats all negative samples equally, potentially misclassifying semantically similar negatives, which limits its adaptability in cross-modal distillation.
@@
(3) COMODO (ours): COMODO extends contrastive learning by distilling the similarity distribution rather than relying solely on instance-wise alignment. Consistent with our gradient analysis in Eq. 6, unlike InfoNCE, which assumes all negatives are equally dissimilar, COMODO preserves the structure of the teacher’s similarity space, leading to more robust and generalizable feature representations. Figure 8 shows that COMODO consistently outperforms L2 and InfoNCE across all datasets, demonstrating the effectiveness of similarity distribution distillation in cross-modal representation learning.
Figure 8. Accuracy of distillation methods across datasets.
4.6.5. Sampling Methods.
To investigate the effectiveness of the FIFO queue in our cross-modal distillation framework, we conduct an ablation study comparing it against random sampling strategy. For a fair comparison, all other settings, including the number of samples used for distillation in each step (i.e., the queue size for FIFO and the number of samples for random selection), are kept identical.
(1) Empirical Results. As shown in Table 8, FIFO yields superior performance across all datasets. This consistent improvement motivates a deeper investigation into the optimization dynamics induced by the sampling method.
Train Sampling A@1 A@3 A@5
Ego4D FIFO 59.13 78.79 85.65
Random 58.47 78.72 85.64
EgoExo4D FIFO 84.92 98.28 99.59
Random 84.68 98.32 99.59
MMEA FIFO 92.48 99.01 99.77
Random 92.25 98.63 99.70
Table 8. Comparison of FIFO and random sampling.
(2) Theoretical Hypothesis. Sampling as an implicit temporal prior on the minimization flow. We posit that the choice of sampling strategy functions as an implicit regularizer, imposing a distinct prior on the optimization process’s minimization flow. The random sampling approach draws a mini-batch of video embeddings randomly and independently from the entire dataset at each iteration. This results in gradient estimations that are largely independent from one step to the next, potentially leading to a high-variance, discontinuous optimization path. In contrast, the FIFO queue introduces a temporal continuity prior. At any two consecutive steps, $t$ and $t+1$, the queue contents $Q_{t}$ and $Q_{t+1}$ are highly correlated, as only the oldest batch of embeddings is replaced by the newest. This ensures that the data distribution used to compute the distillation loss, $P_{v}(i|Q)$, transitions smoothly over time. Consequently, the loss landscape “seen” by the student model evolves gradually, encouraging a more stable and continuous minimization flow.
(a) Ego4D, K=30
(b) EgoExo4D, K=30
(c) Ego4D, K=50
(d) EgoExo4D, K=50
(e) Ego4D, K=100
(f) EgoExo4D, K=100
Figure 9. Comparison on the different moving average window size $K$ for the gradient path continuity for FIFO (blue) and Random (red) sampling.
(3) Validation and Discussion. Our investigation is conceptually motivated by  Charpiat et al. (2007), who showed that modifying the inner product space for gradient computation can introduce priors that yield smoother optimization paths. While their work focuses on a gradient-level prior by altering the gradient, we explore an alternative: shaping the optimization dynamics by structuring the data distribution, which we term a data-level prior. To validate that this prior indeed leads to a more continuous optimization flow, we design an experiment to visualize the path’s continuity. We compute the gradient of $L_{CE}$ with respect to the IMU student model’s parameters $\theta_{x}$ at each iteration $t$, denoted as $\mathbf{g}_{t}=\nabla_{\theta_{x}}L_{CE}^{(t)}$. We then measure the cosine similarity between the current gradient $\mathbf{g}_{t}$ and the moving average of the preceding $K$ gradients, $\bar{\mathbf{g}}_{t-1}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{g}_{t-k}$. This metric quantifies the consistency of the gradient direction over time; a higher cosine similarity indicates a more stable and coherent optimization trajectory.
To ensure the robustness of our findings, we conducted this analysis across different window sizes for the moving average, specifically for $K\in\{30,50,100\}$. A smaller $K$ makes the measure more sensitive to high-frequency changes in gradient direction, while a larger $K$ provides a more smoothed, long-term estimate of the path’s direction.
The results, visualized in Figure 9, provide compelling and consistent evidence supporting our hypothesis. Across both the Ego4D (left column) and EgoExo4D (right column) datasets, the gradient continuity for the FIFO strategy (blue curves) is markedly consistently higher than that of random sampling (red curves). This superiority of FIFO holds true for all tested window sizes of $K$, while the absolute smoothness of the moving average curves naturally increases with a larger $K$ (as seen by the less jagged lines in the bottom row $K=100$ compared to the top $K=30$).
This robustness analysis confirms that the temporal continuity prior induced by the FIFO queue is not a mere artifact of a specific hyperparameter choice but a fundamental mechanism stabilizing the optimization process. The sustained higher cosine similarity indicates that the FIFO queue induces a more continuous and stable minimization flow, where the optimization trajectory is more coherent over time. We posit that this stabilized gradient direction allows the optimizer to converge more efficiently and effectively, preventing unstable updates and ultimately explaining the superior performance of the FIFO strategy observed in our experiments.
5. Discussion and Implications
Our work demonstrates that cross-modal, self-supervised distillation from video to IMU is a highly effective and scalable paradigm for overcoming the long-standing data scarcity problem in wearable HAR. The success of COMODO not only presents a state-of-the-art technical solution but also offers several broader implications for the design and development of future human-centered ubiquitous technologies.
5.1. Implications for Egocentric HAR: A Path Beyond Manual Annotation
A key implication of COMODO is the shift it suggests in the data collection and development lifecycle for wearable HAR systems. The field has long been hampered by the annotation bottleneck, a challenge frequently cited as a primary obstacle to building robust, generalizable real-world (”in-the-wild”) models (Haresamudram et al., 2025). COMODO offers a practical solution to this long-standing dilemma.
First, our framework acts as a conduit to leverage the immense progress and rich semantic knowledge encapsulated in large-scale, pre-trained vision models. The computer vision community has invested enormous effort in creating powerful foundation models (e.g., TimeSformer (Bertasius et al., 2021), VideoMAE (Tong et al., 2022)) trained on vast and diverse video datasets like Kinetics (Carreira and Zisserman, 2017; Caba Heilbron et al., 2015; Goyal et al., 2017; Grauman et al., 2022, 2024). COMODO provides a principled way for the wearable computing community to directly tap into this wealth of knowledge, transferring the nuanced understanding of human actions from the visual domain to the resource-constrained IMU domain, all without requiring a single IMU label. Our strong cross-dataset generalization results (Table 4) are crucial evidence for this; they indicate that the distilled knowledge captures transferable semantic structure of human activities, rather than superficial, dataset-specific correlations.
Second, COMODO is strategically positioned for the next generation of wearable devices. As smart glasses and other multimodal wearables become more prevalent, the collection of naturally occurring, temporally aligned video and IMU data will become common. In this emerging ecosystem, COMODO offers a way to bypass the need for manual IMU annotation entirely. Instead of facing the daunting and often impractical task of labeling hours of abstract IMU signals, developers can directly leverage these paired, unlabeled data streams, training a high-performance, deployment-ready IMU model directly. Furthermore, by breaking the dependency on IMU datasets with limited activity classes (Zhang and Sawchuk, 2012; Anguita et al., 2013; Reiss and Stricker, 2012; Chavarriaga et al., 2013), COMODO opens the door to recognizing a much wider, more personalized ”long tail” of human activities. Current HAR systems are often limited to a dozen or so generic activities like walking or running. However, real human life is composed of a vast array of nuanced, context-specific activities that span, but are not limited to domestic routines, occupational tasks, creative and leisure pursuits, and social interactions (see Section 4.1.1 for activity categories across datasets). It is infeasible to collect labeled IMU data for all such activities. By leveraging video, which is easier to find for niche activities, our approach provides a scalable path towards creating systems that understand the unique facets of an individual’s life, a promising direction of proactive health and wellbeing applications.
5.2. Implications for Ubiquitous and Wearable Systems: Designing for Practicality
The design of ubiquitous and wearable systems is a constant exercise in balancing computational power, user experience, and practical constraints. COMODO contributes directly to this by addressing the critical trade-off between sensing fidelity and deployment feasibility.
First, our framework provides a concrete methodology for decoupling training-time requirements from deployment-time constraints. Continuous wearable video is often untenable due to power consumption and privacy concerns (Bhardwaj et al., 2024; Mujirishvili et al., 2024; Liao et al., 2025; Opaschi and Vatavu, 2020). COMODO addresses these barriers by learning from video offline and operating with IMU online, reconciling semantic richness with the requirements of unobtrusive, human-centered sensing. It allows systems to benefit from the wisdom of the ”powerful but impractical” modality (video) while deploying the ”practical but data-poor” modality (IMU).
Second, the model-agnostic nature of our framework (demonstrated in Table 5) offers significant flexibility for system designers. It suggests that the COMODO paradigm is not tied to a specific architecture but is a general principle. As more powerful video or time-series foundation models emerge, they can be readily integrated as improved teachers or students, allowing for the continuous evolution of wearable intelligence without redesigning the entire system from the ground up.
5.3. Limitation and Future Work
Despite these advantages, certain constraints remain. First, COMODO relies on the availability of synchronized video-IMU data pairs during training. While this dependency currently limits the use of standalone IMU datasets, the growing adoption of next-generation wearable devices (e.g., smart glasses) is expected to facilitate the collection of naturally occurring, temporally aligned video–IMU data. Second, precise temporal alignment between heterogeneous modalities remains a challenge. Nevertheless, our similarity-distribution alignment inherently mitigates this issue by design; by preserving the teacher model’s semantic structure, COMODO ensures that semantically similar segments are not equally penalized as negatives. Lastly, COMODO’s effectiveness is inherently tied to the representational capacity of the pretrained video teacher, although its model-agnostic design allows it to benefit from the continuous advancement of video foundation models. Future research could focus on enhancing robustness against more severe asynchronicity to further reduce data requirements.
While our current framework employs a one-way knowledge transfer from a pretrained, frozen teacher, future work could investigate mutual distillation frameworks, where the video and IMU models learn collaboratively, potentially allowing the teacher to also refine its representations based on the fine-grained temporal patterns captured by the IMU. Additionally, our current setup does not explicitly consider personalization or contextual adaptation; integrating continual learning (Parisi et al., 2019; Leite and Xiao, 2022) and user-specific adaptation mechanisms could make the framework more responsive to individual behavioral nuances. Finally, the distillation paradigm proposed in COMODO is not limited to HAR. We believe this approach of leveraging a data-rich modality to supervise a data-scarce modality may serve as a useful paradigm for other ubiquitous sensing tasks, such as transferring knowledge from video to audio for acoustic scene analysis or from video to physiological sensors for affective computing. We hope this perspective can inspire future work toward more general and flexible multimodal learning frameworks for ubiquitous and human-centered AI.
6. Conclusion
In this work, we introduced COMODO, a cross-modal self-supervised distillation framework designed to address a fundamental trade-off in human-centered computing: the interplay between the rich semantics of video and the practical deployability of IMU sensors. By distilling knowledge from a powerful, frozen video encoder to a lightweight IMU encoder via a dynamic instance queue, COMODO effectively bypasses the need for costly manual annotation of sensor data. Our extensive experiments demonstrate that this approach not only matches or surpasses supervised baselines but, more importantly, learns representations with strong cross-dataset generalization, a critical requirement for real-world utility. COMODO’s model-agnostic and flexible design offers a principled and scalable path forward, enabling future wearable systems to leverage the immense progress in visual foundation models. We believe this work represents a significant step towards creating more capable, efficient, and privacy-preserving wearable AI.
Acknowledgements.